OpenStudy (anonymous):

Find the equation of the axis of symmetry of y=4x² - 2x - 1.

OpenStudy (anonymous):



OpenStudy (campbell_st):

axis if symmetry is at x = -b/2a in your question a = 4 and b = -2

OpenStudy (anonymous):

so the axis of symmetry is -.25?

OpenStudy (anonymous):

well x=-.25 ?

OpenStudy (campbell_st):

no its -(-2)/2x4 = 0.25

OpenStudy (anonymous):

oh okay i distributed the negative my mistake. thanks!